The stock price of XYZ Ltd is $45. EPS is $2.50 and the required return on equity is 8 per cent annually. Which one of the following statements is false?

The answer is - "All else equal, an increase in the required return on equity to 10 per cent will cause NPVGO to rise to $22 per share"

But I don't know how it arrived to that answer.
Answer #1

PVGO = Price - Earning/Cost of Equity =45-2.50/8% =13.75

New PVGO =45-2.5/10% =20

NPVGO is 20 per share . and not 22 .